OpenStudy (anonymous):

Solve 5n+20n=5(n+20)

OpenStudy (anonymous):

5n+20n=5(n+20) You have to isolate 'n' by moving everything else to the other side of the equal sign. 1. Right Side - Multiply the '5' into the brackets. 5n+20n=5n+100 2. Left Side - Add like terms. 25n=5n+100 3. Right Side - Move '5n' to the left side, switching the signs on it. (-5n) Subtract like terms. 25n-5n=100 20n=100 4. To isolate the 'n' by itself, divide the '20' from '20n'. What you do to one side, you do to the other, so also divide '20' from '100'. (20n)/20=100/20 n=100/20 =5 n=5

OpenStudy (anonymous):

Thats whati got! Thanks!!!